OpenStudy (lolacole12):

what is the value of the expression x - y if x=4.51 y=-2.97

OpenStudy (lolacole12):

HI ill give medal

Vocaloid (vocaloid):

it tells you that x is equal to 4.51 and y is equal to -2.97 so what is x minus y?

OpenStudy (anonymous):

4.51 - -2.97

OpenStudy (lolacole12):

options is a.-2.54 B.-1.54 C. 6.48 D.7.48

OpenStudy (lolacole12):

@Vocaloid

OpenStudy (lolacole12):

@tyler.ford

Vocaloid (vocaloid):

4.51 - (-2.97) = ?

OpenStudy (lolacole12):

Its - 1.54

OpenStudy (lolacole12):

right @Vocaloid

Vocaloid (vocaloid):

nope. 4.51 - (-2.97) = 4.51 + 2.97 = ?

OpenStudy (anonymous):

Simple problem...Haha...

OpenStudy (lolacole12):

so C

Vocaloid (vocaloid):

nope, not C

OpenStudy (lolacole12):

no d

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

LOL

OpenStudy (lolacole12):

one more

OpenStudy (anonymous):

Sorry, It's pretty funny. xD

OpenStudy (lolacole12):

what value of (a) makes the equation true? a+10.8=-7.5 A. -18.3 C. 3.3 B.-3.3 D. 17.3 @Vocaloid

Vocaloid (vocaloid):

subtract 10.8 from both sides

OpenStudy (lolacole12):

C

OpenStudy (lolacole12):

@Vocaloid